Simplify: -4-[3-4(2x-5)]
Answer by nyc_function(2741)

-4-[3-4(2x-5)] =

-4 -[3 - 6x + 20] =

-4 -[-6x + 23] =

-4 + 6x - 23 =

-27 = -6x

-27/-6 = x


27/6 = x


9/2 = x